# Theoretical limits of magnetic detection of structural surface defects at the nanometer scale

### v1

@@ Abstract
We present a theoretical study on the magnetic signals of structural surface defects like cracks or indents combined with rough surfaces or subsurface inclusions of soft ferromagnetic metals like body-centered cubic Fe or amorphous CoFeB. We discuss limits of early detection of small surface defects on the basis of calculated magnetic stray fields few tens of nm above the surface. The considered surface imperfections have extensions of a few nm which correspond to low multiples of the magnetic exchange lengths of Fe or CoFeB. The detection of such small inhomogeneities requires that the sensor is about as close to the surface as the size of the inhomogeneity is. Furthermore, the step width of a scanning sensor must be of the same size as well. Both these requirements may be fulfilled for instance by scanning microscopy with diamond nitrogen-vacancy-center quantum sensors.
I Introduction
@@
Detection of structural surface imperfections like few nm small cracks or indents is of high technical relevance for monitoring and estimating the time to failure of structural materials. Failure of metals by fatigue or fracture often starts at the surface of a piece of metal. Non-destructive testing (NDT) is an established approach that can reduce maintenance costs and help to improve reliability, durability and safety of components. There are numerous macro-scale techniques like visual inspection, resistance strain gauges, piezoelectric transducers, vibrational and modal analysis, as well as electric and magnetic sensors.[1]
@@
The purpose of NDT is to detect the local origin of failure as early as possible, and it is not limited to the macro-scale. For example, at the micro-scale, cracks with an opening of only 5 m can be detected[3, 4] with the near-field microwave technique.[2] The next step is consequently to address the nanometer-scale, which is the focus of this work. Magnetic sensors like scanning superconducting microscopy (SSM) based on SQUIDs (Superconducting QUantum Interferometer Devices), magnetic force microscopy (MFM) and scanning nitrogen-vacancy (NV) centre microscopy (SNVM) have currently the highest spatial resolution down to 10m. Furthermore, the latter three mentioned methods are sensitive to small fields down to T (for more information see e.g. the recent review of Marchiori et al.[5] and references therein) which is sensitive enough for nm-sized defects as we will show in the following.
@@
However, real materials have rough surfaces and do not contain only indents or cracks but also subsurface inclusions or surface protrutions. Thus, the magnetic stray field signal above the surface of the sample under investigation may be hard to interpret. With our theoretical study we want to give answers to several questions which can help to optimize the techniques and to clarify the requirements for such measurements. First of all: how strong are the signals of nm-sized surface defects like cracks or indents? Furthermore: how does the signal of an indent scale with its volume and depend on its shape? What is the influence of a rim around an indent? What is the effect of other surface imperfections nearby? How strong is the effect of a magnetically different subsurface inclusion below an indent? And finally: how does the magnetic signal depend on the axial distance between the metal surface and the sensor, and how dense does the lateral step width of the scanning sensor need to be?
@@ II.1 Simulation Details
Since the computational cost of MuMax3 is dominated by fast Fourier transform (FFT) using the cuFFT library, which performs best for sizes of powers of two[6] we have used discretized simulation boxes of either 512512256 nm (large size) or 512512128 nm (small size).
To achieve a numerically well enough converged total energy, angles between discrete spins (normalized magnetization vectors) in adjacent cells should be less than about 30° by choosing a small enough size for the mesh cells.[6] With a size of nm for both CoFeB and Fe this angle criterion was well fulfilled.
@@ II.2 Model systems
3. Indent with rim and additional imperfections nearby on a rough surface: The surface roughness was modeled by arbitrarily putting five little cubes onto the flat surface. These cubes have an edge length of 8 nm (see Fig. 6a).
4. Indent on a rough surface above a subsurface inclusion: As an additional feature we modeled subsurface inclusions by spheres below the indents with reduced saturation magnetization of 50%, 80% or 90% with respect to the bulk material value . This case is a simple model for a small precipitate (e.g., a cementite particle) or a small plastically phase-transformed region (e.g., a fcc or hcp iron particle) below an indent on a surface of e.g. the bcc iron, which has slightly different magnetic properties than iron.
@@ III.1 Signals of indents
Moving from the cubic indent to cuboidal indents with different edge lengths in x and y directions, one gets magnetic field signals that differ significantly in x and y directions. The lowest-energy solution is always the one where the magnetization is pointing perpendicular to the smallest side wall of the indent: e.g. if L L then A= LL A= LL, and the magnetization points in [100] direction. If edge lengths of the indent are similar in all directions, like in the cases of cubic, hemispherical and pyramidal indents, then the size and the shape of the signal nearly do not depend on the direction parallel to the surface plane. A typical signal produced by a hemispherical indent of diameter d= 24 nm is plotted in Fig. 3 (blue line). At 50 nm distance of the sensor above the surface one has a stray field signal of about 0.7 mT and the signal width in x or y directions is about 60 nm (we have taken the width at decay to 1/e).
@@
![Refer to caption](extracted/5467832/Fig3.png)
@@ III.2 Signals of indents with rims
The indent can only be resolved as a dip in the magnetic stray-field signal by approaching the sensor closer to the surface. This is shown for the hemispherical indent in Fig. 3. At a distance of 40 nm between sensor and surface (orange line in Fig. 3) one resolves two peaks coming from the rim. Approaching the sensor to 30 nm the signal is even more pronounced: the dip in the middle caused by the indent is superposed with two peaks at the left and right side due to the rim.
@@
For larger distances beween sensor and surface the difference goes to zero. Already at nm, about twice the diameter of the hemispherical indent, the dip and the pair of peaks cannot be resolved any more (compare the green lines in Fig 5a) with the red line in Fig. 3).
@@
From the experimental perspective the positioning of a sensor to such a short distance above a metal surface is a very ambitious task. To our knowledge the best signal resolutions were achieved by scanning-probe-microscopy set-ups for distances between 15 and 25 nm with SNVM.[15, 16] The advantage of the SNVM over the SSM and MFM is the potentially very small sensor size since the single NV-center defect complex in diamond employed as the sensor is of an atomic size. NV-center sensors, which have sufficiently long coherence times for a measuring process, can be implanted approximately 10 nm below the diamond surface[17] and the diamond surface of the sensor can in principle be put in contact with the surface of the soft-magnetic metal specimen, which means that a spatial sensor resolution of 10 nm is feasible.
@@ III.3 Effect of additional imperfections on the surface
An important complication of detecting inhomogeneities in or on a surface is a hardly avoidable surface roughness. As a simple structural model for this we add five tiny cubes of edge length 8 nm (called particles in the following) on the otherwise flat surface with the indent. We assume that these cubes are not too near to the indent, i.e. within a radius of a few nm, as this arrangement would have a similar effect as the previously discussed rim on the magnetic signal of the indent.
However, as can be see from Fig. 6, even if the cubes are rather far away from the indent the signal becomes strongly disturbed and the hemispherical indent, which has an approximately seven times larger volume than the other few particles on the surface (an indent with diameter d= 24 nm), can only be seen when the sensor directly passes above the center of the indent (compare red and orange curves for flat and slightly rough surfaces). By passing the indent with the sensor only 50 nm away in lateral direction leads to a too small signal which is covered by the background variation coming from the other few small particles (compare green and blue curves with orange curve). One has to increase the size of a hemispherical indent to a diameter nm, which corresponds to a volume that is 28 times bigger than that of the other small particles, to get a significant stray-field signal that exceeds the background variation of about 400 T due to the few other small particles in our model (compare green and black curves).
Hence, sensing magnetic fields above macroscopically smooth surfaces, which are microscopically rough due to arbitrarily dispersed, nm-sized particles, require a dense scanning grid. Since surface defects like indents or cracks can have other small surface inhomogeneities nearby, we conclude that the scanning grid needs to have a step size which is of order of the dimension of the surface defect. This may be technically cumbersome as it results in long times needed for scanning surface areas of m sizes. Furthermore, based on our simulation results we suggest that indents need to have at least about 30 times larger volume than the other objects on the surface for being reliably detectable by a scanning magnetic-field sensor of such high spatial resolution and high signal sensitivity as, e.g., a modern SNVM.[15, 16]
@@ IV Summary
A significant complication for the interpretation of magnetic stray-field signals of the surface defects of interest (indents or cracks) comes from other small and randomly dispersed imperfections on a macroscopically smooth, microscopically rough metal surface. Our micromagnetic simulation results illustrate that such an atomic-scale surface roughness can cover the signal of indents since their magnetic signals are decaying quickly over few tenth of nm. This means that a sensor that bypasses the center of a nm-sized indent by only about 50 nm will not give a clear dip or peak in the magnetic stray-field signal that is distinguishable from a background variation unless the indent is more than at least 30 times bigger in volume than other defect objects above or below the surface. This implies the need of a dense scanning grid for the magnetic sensor to detect surface defects of few nm size reliably. However, practically this may require long measuring times with stable scanning-sensor set-ups to get magnetic stray-field maps of micrometer-sized areas of magnetic metal surfaces.

### v2

@@ Abstract
We present a theoretical study on the magnetic signals of structural surface defects like cracks or indents combined with inhomogeneities on the surface or subsurface inclusions of soft ferromagnetic metals like body-centered cubic Fe or amorphous CoFeB. We discuss limits of early detection of small surface defects on the basis of calculated magnetic stray fields few tens of nm above the surface. The considered surface imperfections have extensions of a few nm which correspond to low multiples of the magnetic exchange lengths of Fe or CoFeB. The detection of such small inhomogeneities requires that the sensor is about as close to the surface as the size of the inhomogeneity is. Furthermore, the step width of a scanning sensor must be of the same size as well. Both these requirements may be fulfilled for instance by scanning microscopy with diamond nitrogen-vacancy-center quantum sensors.
pacs:
Keywords: crack detection, magnetic field measurment, NV sensorI Introduction
@@
Detection of structural surface imperfections like few nm small cracks or indents is of high technical relevance for monitoring and estimating the time to failure of structural materials. Failure of metals by fatigue or fracture often starts at the surface of a piece of metal. Non-destructive testing (NDT) is an established approach that can reduce maintenance costs and help to improve reliability, durability and safety of components. There are numerous macro-scale techniques like visual inspection, resistance strain gauges, piezoelectric transducers, vibrational and modal analysis, as well as electric and magnetic sensors.[1, 2]
For example, at the micro-scale, cracks with an opening of only 5 m can be detected [3, 4] with the near-field microwave technique.[5] The next step is consequently to address the nanometer scale, which is the focus of this work. To detect the origins of damage and failure as early and as locally as possible, for example at a weld seam of a steely (hydrogen) gas pipeline, or other metallic building components which are exposed to static or cyclic loading, is advantageous in various aspects. The observation and tracking of nanometer-sized defects can improve component-design or lifetime-assessment models with respect to mechanical fracture or fatigue. This is highly desirable since current codes and standards dealing with damage and failure of metals, need to be adapted to metals in heavy-duty environments [6], e.g. in hydrogen-gas atmospheres.[7, 8] An actual topic is the investigation of hydrogen related local damage in microstructures of metals. This has been studied recently in e.g. Nickel oligocrystals by Singh et al.[9]. The observation of initiation and propagation of microstructural damage in such macroscale specimens is of importance for improvement of understanding. However, the detection of material damage from the beginning at the nanometer scale will be even better [10, 11, 12, 13]. With our micromagnetic model calculations we estimate magnitude and spatial distribution of magnetic stray fields due to nanometer-sized indents and further inhomogeneities on magnetic metal surfaces which is useful for the design and layout of magnetic quantum sensors for NDT of materials.
@@
Magnetic sensors like scanning superconducting microscopy (SSM) based on SQUIDs (Superconducting QUantum Interferometer Devices), magnetic force microscopy (MFM) and scanning nitrogen-vacancy (NV) centre microscopy (SNVM) have currently the highest spatial resolution down to 10-8m. Furthermore, the latter three mentioned methods are sensitive to small fields down to 1 T (for more information see e.g. the recent review of Marchiori et al.[14] and references therein) which is sensitive enough for nm-sized defects as we will show in the following.
@@
However, real materials have surfaces and do not contain only indents or cracks but also subsurface inclusions or surface protrutions. Thus, the magnetic stray field signal above the surface of the sample under investigation may be hard to interpret. With our theoretical study we want to give answers to several questions which can help to optimize the techniques and to clarify the requirements for such measurements. First of all: how strong are the signals of nm-sized surface defects like cracks or indents? Furthermore: how does the signal of an indent scale with its volume and depend on its shape? What is the influence of a rim around an indent? What is the effect of other surface imperfections nearby? How strong is the effect of a magnetically different subsurface inclusion below an indent? And finally: how does the magnetic signal depend on the axial distance between the metal surface and the sensor, and how dense does the lateral step width of the scanning sensor need to be?
@@ II.1 Simulation Details
Since the computational cost of MuMax3 is dominated by fast Fourier transform (FFT) using the cuFFT library, which performs best for sizes of powers of two[15] we have used discretized simulation boxes of either 512512256 nm3 (large size) or 512512128 nm3 (small size).
To achieve a numerically well enough converged total energy, angles between discrete spins (normalized magnetization vectors) in adjacent cells should be less than about 30° by choosing a small enough size for the mesh cells.[15] With a size of nm3 for both CoFeB and Fe this angle criterion was well fulfilled.
@@ II.2 Model systems
3. Indent with rim and additional imperfections nearby on a surface: The inhomogeneities on the surface were modeled by arbitrarily putting five little cubes onto the flat surface. These cubes have an edge length of 8 nm (see Fig. 6a).
4. Indent and inhomogeneit on a surface above a subsurface inclusion: As an additional feature we modeled subsurface inclusions by spheres below the indents with reduced saturation magnetization of 50%, 80% or 90% with respect to the bulk material value . This case is a simple model for a small precipitate (e.g., a cementite particle) or a small plastically phase-transformed region (e.g., a fcc or hcp iron particle) below an indent on a surface of e.g. the bcc iron, which has slightly different magnetic properties than iron.
@@ III.1 Signals of indents
Moving from the cubic indent to cuboidal indents with different edge lengths in x and y directions, one gets magnetic field signals that differ significantly in x and y directions. The lowest-energy solution is always the one where the magnetization is pointing perpendicular to the smallest side wall of the indent: e.g. if Lx Ly then Ax= LxLz Ay= LyLz, and the magnetization points in [100] direction. If edge lengths of the indent are similar in all directions, like in the cases of cubic, hemispherical and pyramidal indents, then the size and the shape of the signal nearly do not depend on the direction parallel to the surface plane. A typical signal produced by a hemispherical indent of diameter d= 24 nm is plotted in Fig. 3a (blue line). At 50 nm distance of the sensor above the surface one has a stray field signal of about 0.7 mT and the signal width in x or y directions is about 60 nm (we have taken the width at decay to 1/e).
@@ III.2 Signals of indents with rims
The indent can only be resolved as a dip in the magnetic stray-field signal by approaching the sensor closer to the surface. This is shown for the hemispherical indent in Fig. 3a. At a distance of 40 nm between sensor and surface (orange line in Fig. 3a) one resolves two peaks coming from the rim. Approaching the sensor to 30 nm the signal is even more pronounced: the dip in the middle caused by the indent is superposed with two peaks at the left and right side due to the rim.
@@
For larger distances beween sensor and surface the difference goes to zero. Already at nm, about twice the diameter of the hemispherical indent, the dip and the pair of peaks cannot be resolved any more (compare the green lines in Fig 5a) with the red line in Fig. 3a).
@@
From the experimental perspective the positioning of a sensor to such a short distance above a metal surface is a very ambitious task. To our knowledge the best signal resolutions were achieved by scanning-probe-microscopy set-ups for distances between 15 and 25 nm with SNVM.[24, 25] The advantage of the SNVM over the SSM and MFM is the potentially very small sensor size since the single NV-center defect complex in diamond employed as the sensor is of an atomic size. The SQUID’s loop and the size of the MFM sensor are typically in the range of micrometers.[26]
NV-center sensors, which have sufficiently long coherence times for a measuring process, can be implanted approximately 10 nm below the diamond surface[27] and the diamond surface of the sensor can in principle be put in contact with the surface of the soft-magnetic metal specimen, which means that a spatial sensor resolution of 10 nm is feasible.
@@ III.3 Effect of additional imperfections on the surface
An important complication of detecting indents or cracks on a surface is a hardly avoidable surface imperfection. As a simple structural model for this we add five tiny cubes of edge length 8 nm (called particles in the following) on the otherwise flat surface with the indent. We assume that these cubes are not too near to the indent, i.e. within a radius of a few nm, as this arrangement would have a similar effect as the previously discussed rim on the magnetic signal of the indent.
However, as can be see from Fig. 6, even if the cubes are rather far away from the indent the signal becomes strongly disturbed and the hemispherical indent, which has an approximately seven times larger volume than the other few particles on the surface (an indent with diameter d= 24 nm), can only be seen when the sensor directly passes above the center of the indent (compare red and orange curves for flat and slightly imperfect surfaces). By passing the indent with the sensor only 50 nm away in lateral direction leads to a too small signal which is covered by the background variation coming from the other few small particles (compare green and blue curves with orange curve). One has to increase the size of a hemispherical indent to a diameter nm, which corresponds to a volume that is 28 times bigger than that of the other small particles, to get a significant stray-field signal that exceeds the background variation of about 400 T due to the few other small particles in our model (compare green and black curves).
Hence, sensing magnetic fields above macroscopically smooth surfaces, which are microscopically inhomogeneous due to arbitrarily dispersed, nm-sized particles, require a dense scanning grid. Since surface defects like indents or cracks can have other small surface inhomogeneities nearby, we conclude that the scanning grid needs to have a step size which is of order of the dimension of the surface defect. This may be technically cumbersome as it results in long times needed for scanning surface areas of m2 sizes. Furthermore, based on our simulation results we suggest that indents need to have at least about 30 times larger volume than the other objects on the surface for being reliably detectable by a scanning magnetic-field sensor of such high spatial resolution and high signal sensitivity as, e.g., a modern SNVM.[24, 25]
@@ IV Summary
A significant complication for the interpretation of magnetic stray-field signals of the surface defects of interest (indents or cracks) comes from other small and randomly dispersed imperfections on a macroscopically smooth surface. Our micromagnetic simulation results illustrate that such atomic-scale surface inhomogeneities can cover the signal of indents since their magnetic signals are decaying quickly over few tenth of nm. This means that a sensor that bypasses the center of a nm-sized indent by only about 50 nm will not give a clear dip or peak in the magnetic stray-field signal that is distinguishable from a background variation unless the indent is more than at least 30 times bigger in volume than other defect objects above or below the surface. This implies the need of a dense scanning grid for the magnetic sensor to detect surface defects of few nm size reliably. However, practically this may require long measuring times with stable scanning-sensor set-ups to get magnetic stray-field maps of micrometer-sized areas of magnetic metal surfaces.